OpenStudy (anonymous):

find the product to 1 1/3 and 5/6

OpenStudy (aakashsudhakar):

So you're multiplying 1&(1/3) and (5/6) together? 1 & 1/3 becomes 4/3 to make it simpler. (4/3) x (5/6) = 20/18 = 10/9

OpenStudy (anonymous):

ok thank you you are a life saver